can you please help me with this
john rents a car for $85 a week plus $0.26per mile.how far can he travel on a budget of $215

thanks alot

Answer by ankor@dixie-net.com(22740)

John rents a car for $85 a week plus $0.26per mile.how far can he travel on a budget of $215
:
Let d = miles travels
:
.26d + 85 = 215
.26d = 215 - 85
.26d = 130
d = 130/.26
d = 500 mi, he can travel on $215
:
:
Check solution in original equation
.26(500) + 85
130 + 85 = 215